A person having the nearest distance of distinct vision of 32 cm uses a reading lens of 8 cm focal length . The magnification of his reading lens is 5 4 3 2

  • A person having the nearest distance of distinct vision of 32 cm uses a reading lens of 8 cm focal length . The magnification of his reading lens is
  • 5
  • 4
  • 3
  • 2

Grade:8

1 Answers

Saurabh Koranglekar
4 years ago
We know that,

magnification of lens is m = 1 + D/f

So, m = 1+ 32/8

= 1 + 4

Thus, m = 5